If a sphere is inscribed in a cube of side What is the volume of the sphere?
1
Expert's answer
2020-12-14T12:45:45-0500

If the sphere is inscribed in a cube of side "a" , then its diameter is "d=a."

Then the radius of the sphere is "r=\dfrac{d}{2}=\dfrac{a}{2}."

 The volume of the sphere is


"V=\dfrac{4}{3}\pi r^3=\dfrac{4}{3}\pi (\dfrac{a}{2})^3"

"=\dfrac{\pi a^3}{6} (units^3)"

"V=\dfrac{\pi a^3}{6} cubic \ units"